if ray QS bisects angle RQT, and if angle RQS=7x+7 and angle RQT=9x+54, find angle SQT.
Answer by mananth(16949)

if ray QS bisects angle RQT, and if angle RQS=7x+7 and angle RQT=9x+54, find angle SQT.
Angle RQS + angle SQT = angle RQT
Angle SQT = Angle RQT - Angle RQS = Angle SQT
9x+54 -(7x+7)= Angle SQT
9x +54 -7x-7 = Angle SQT
2x +47 = Angle SQT
Since ray QS bisects angle RQT Angle RQS = angle SQT
7x+7 = 2x+47
5x = 40
x =8
Angle SQT = 2x+47 =63 deg